solve for x
5ln2x-7=11

Answer by stanbon(75887)

solve for x
5ln(2x-7) = 11
-----
ln(2x-7) = 11/5 = 2.2
----
2x-7 = e^2.2
---
2x = 7+e^2.2
-----
x = (7+e^2.2)/2
---
x = 8.0125
===============
Cheers,
Stan H.
===============